In Δ​XYZ, measure of angle∠Y equals=78​, XY equals=99​, and XZ equals=13. To the nearest​ tenth, what is measure of angle∠​Z? This is apart of the law sines. Thanks.
Answer by Boreal(15235)

This has to be drawn. I am assuming that XY= 9.9, otherwise this can't be solved.
y/sin Y=z/sin Z
13/sin 78=99/sin Z
13.29=9.9/sin Z
sin Z=9.9/13.29
=0.7449.
take the arc sin of that, and Z=48.15 degrees.